# Effortless, Simulation-Efficient Bayesian Inference using Tabular Foundation Models

### v1

@@ 3 Experiments
To assess NPE-PF, we conduct experiments on synthetic SBI benchmark tasks and real data, covering scenarios from low to high-dimensional data and including cases with model misspecification.
Figure 2: SBI benchmark results for amortized and sequential NPE-PF. (a) C2ST for NPE, NLE, and NPE-PF across ten reference posteriors (lower is better); dots indicate averages and bars show 95% confidence intervals over five independent runs. (b) Average time to generate $10^{4}$ posterior samples and (if applicable) training time. (c) C2ST for sequential methods.
@@ 3.1 Amortized and sequential benchmark performance
We evaluate NPE-PF on various tasks from the SBI benchmark [24], which provides ground truth posterior samples for 10 observations for each task. We measure posterior sample quality using the classifier two-sample test (C2ST, 48). A C2ST accuracy of 0.5 indicates that the approximate posterior exactly matches the ground truth posterior, as the classifier fails to distinguish between the two sample sets. In contrast, an accuracy of 1.0 reflects a strong mismatch between the estimated and true posterior. As baselines, we compare against flow-based NPE and NLE provided by the SBI library [49, 50]. For sequential baselines, we use truncated sequential NPE (TSNPE [34]) and sequential NLE (SNLE [11]). Posterior estimates are computed across four simulation budgets, ranging from $10^{2}$ to $10^{5}$ simulations. Beyond $10^{4}$ simulations, NPE-PF incorporates filtering, whereas an alternative unfiltered variant, denoted NPE-PF (unfiltered), extends the context size beyond the TabPFN-recommended limit. For budgets $\leq 10^{4}$, both variants are identical.
@@
Amortized inference. For $10^{2}$ simulations, NPE-PF substantially outperforms NLE and NPE on most tasks (Fig. 2a,b). With $10^{3}$ or $10^{4}$ simulations, NPE-PF is generally competitive, though some baselines outperform it on some tasks. For example, NLE outperforms NPE-PF on the simple-likelihood-complex-posterior (SLCP) task, as NPE-PF, like NPE, directly targets the posterior. However, NPE-PF performs substantially better than both baselines on tasks such as SIR or Lotka-Volterra. With $10^{5}$ simulations, the difference between NPE-PF and its unfiltered variant becomes evident. The extended context does not improve performance for larger simulation budgets, while the filtered NPE-PF variant achieves significantly better C2ST accuracies, matching or surpassing the baselines. This advantage also remains in comparison to NRE or NPE ensembles (Appendix Sec. D.2, Fig. D-2), and NPE-PF posterior estimates are well-calibrated (Fig. D-3). NPE-PF is slower than NPE, and comparable to NLE in inference time (Fig. 2b). In cases where data dimension $d_{x}$ outgrows the supported feature dimension and learned summary statistics are required, the end-to-end trained NPE, however, outperforms NPE-PF with independently pre-trained summary statistics (Fig. D-4, details in Appendix Sec. D.3).
@@
Sequential inference. We adapt the setup in Deistler et al. [34] to evaluate TSNPE-PF. Rejection sampling is performed using the fast, approximate ratio-based densities (Sec. 2.4). We equally divide the simulation budget into 10 rounds, with proposals truncated to the $1-\varepsilon$ highest density region, where $\varepsilon=10^{-3}$. TSNPE-PF excels with small simulation budgets (Fig. 2c). On the SIR task, it reaches close to optimal C2ST accuracy with only $10^{2}$ simulations. At larger budgets, TSNPE-PF maintains a strong performance, particularly on the Lokta-Volterra task, which is highly challenging for other SBI methods. TSNPE-PF benefits from NPE-PF’s ability to generate high-quality posterior estimates with few simulations, allowing it to acquire more relevant simulations in early rounds.
In summary, NPE-PF provides accurate posterior estimates, especially for small simulation budgets. With filtering, NPE-PF is also competitive for larger simulation budgets. As an in-context learning method, it requires no training on specific simulations and is similar in inference time to other baseline methods. This flexibility makes NPE-PF applicable to a wide range of simulators and setups. In addition to posterior estimation with NPE-PF, we also evaluate the unconditional high-dimensional density estimation capabilities of TabPFN on UCI datasets [51], showing that is more data-efficient than a neural spline flow (Appendix Sec. D.1, Fig. D-1), and perform comparisons between autoregressive and ratio-based density evaluation (Appendix Sec. D.4, Fig. D-5).
@@
Figure 3: Robustness under misspecification. C2ST against a well-specified ground-truth posterior. The red star marks the well-specified model. Rows: Prior and likelihood misspecification. Columns: NPE, NPE-PF, and their difference in terms of C2ST accuracy (darker blue indicates that NPE-PF is better).
@@ 3.2 Impact of misspecification
To investigate how NPE-PF handles misspecification, we use the 2D Gaussian means misspecification benchmark from Schmitt et al. [52], considering both prior and likelihood misspecification. The ground truth is given by the prior ${\bm{\mu}}\sim\mathcal{N}({\bm{0}},{\bm{I}})$ and likelihood ${\bm{x}}_{i}\sim\mathcal{N}({\bm{\mu}},{\bm{I}})$. Prior misspecification uses the prior ${\bm{\mu}}\sim\mathcal{N}({\bm{\mu}}_{m},\tau_{m}{\bm{I}})$ while keeping the true likelihood; Likelihood misspecification uses the true prior with the likelihood ${\bm{x}}_{i}\sim\lambda_{m}\text{Beta}(2,5)+(1-\lambda_{m})\mathcal{N}({\bm{\mu}},\tau_{m}{\bm{I}})$, where $\tau_{m}\in\mathbb{R}^{+}$, $\lambda_{m}\in[0,1]$. We evaluate the C2ST of predicted posteriors against the reference posterior distribution of the well-specified data using $10^{2}$ simulations (details in Appendix Sec. 3.2).
Both NPE-PF and NPE degrade as the level of misspecification increases, with prior misspecification leading to more severe performance degradation than likelihood misspecification (Fig. 3). However, NPE-PF achieves better C2ST values than NPE across a broader range of misspecification levels. In addition, NPE-PF remains insensitive to various feature and noise distributions (Appendix Sec. D.5, Fig. D-6) as well as the choice of autoregressive sample order (Appendix Sec. D.6, Fig. D-7)
Thus, the trend observed in the SBI benchmark extends to these simple, but misspecified tasks, on which NPE-PF also consistently outperforms NPE. Next, we evaluated whether this also holds true on scientific tasks with real data, on which the simulator is potentially misspecified.
Figure 4: Posterior inference for observations from the Allen cell type database. (a) Average distance to observation in standardized space of summary statistics for both the real and synthetic observations. (b) Posterior predictive simulations for real observations using TSNPE-PF and TSNPE. (c) Posterior marginals for one synthetic observation; TSNPE-PF marginals are substantially more constrained for several parameters. (d) Simulation-based calibration for NPE-PF and NPE on the HH simulator.
@@ 3.3 Single-Compartment Hodgkin-Huxley Model
We evaluate the performance of (TS)NPE-PF on a challenging inference task, considering the Hodgkin-Huxley (HH) simulator of single-neural voltage dynamics [53, 54]. Following Gonçalves et al. [55], the simulator has eight parameters and seven summary statistics based on the voltage trace. We infer posteriors for 10 real observations from the Allen cell type database [56], for which inference is known to be challenging [57, 58, 59]. We evaluate inference quality via posterior predictives and the average Euclidean distance between predictives and real observations in a standardized summary statistics space. In addition, for each real observation, we create an associated synthetic and therefore well-specified observation, where the ground truth parameter is known, allowing for direct comparison between inferred posteriors and true values.
TSNPE-PF yields posterior predictives that closely match real observations with a simulation budget of $10^{4}$ (Fig. 4a,b). It outperforms TSNPE, which struggles to produce realistic predictives, especially at smaller budgets. For synthetic observations, TSNPE-PF again achieves better predictive accuracy with fewer simulations, with the average distance approaching zero for a simulation budget of $10^{4}$. In parameter space, TSNPE-PF posteriors are tightly concentrated around the true parameters, with significant improvement over TSNPE (Fig. 4c). In the amortized setting, NPE-PF shows good calibration for $10^{3}$ and $10^{5}$ simulations but is slightly overconfident for $10^{4}$ (Fig. 4d). However, (TS)NPE-PF achieves significantly better prediction quality and yields tight posteriors around the true parameters, making it more sensitive to miscalibration.
These results demonstrate TSNPE-PF’s superior simulation efficiency, achieving the same predictive quality as TSNPE with an order of magnitude fewer simulations. This makes TSNPE-PF particularly well-suited for inference with real, potentially misspecified simulators under limited simulation budgets.
@@ 3.4 Pyloric Network Model
Finally, we evaluate TSNPE-PF on a high-dimensional simulator of the pyloric network in the stomatogastric ganglion (STG) of the crab Cancer Borealis, a well-studied circuit generating rhythmic activity [60, 61, 55]. This model consists of three neurons, each with eight membrane conductances, interconnected by seven synapses, resulting in a 31-dimensional parameter space. The simulated voltage traces are summarized using 15 established statistics. A key challenge in this setting is the extreme sparsity of valid simulations. Specifically, 99% of parameter samples from the prior yield implausible voltage traces, preventing the computation of summary statistics [62].
For this reason, earlier work relied on millions of simulations for an amortized NPE-based posterior approximation (18 million in Gonçalves et al. [55]). This could later be reduced by restricting the prior to valid simulation using an additional classifier [62]. In addition, several sequential algorithms have been developed to tackle this problem [18, 34, 63], the latest of which achieves good posterior predictives using $1.5\cdot 10^{5}$ simulations.
Figure 5: Results on the pyloric simulator. (a) Voltage traces from the experimental measurement (top) and a posterior predictive simulated using the posterior mean from TSNPE-PF as the parameter (bottom). (b) Average distance (energy scoring rule) to observation and percentage of valid simulation from posterior samples; compared to experimental results obtained in Glaser et al. [63].
To perform sequential inference on the pyloric network with TSNPE-PF, we use a restricted prior based on the TabPFN classifier. Due to the high dimensionality of the parameter space, rejection sampling is prohibitive. Therefore, we use sampling importance resampling with an oversampling factor of 10. In the first round, we use 5000 simulations and run TSNPE-PF for 45 rounds, adding 1000 simulations per round for a total of $5\cdot 10^{4}$ simulations (details in Appendix Sec. C.4).
Generated voltage traces from posterior mean estimates of TSNPE-PF closely resemble true measurements (Fig. 5). With more than $50\%$ valid simulations at just $5000$ simulations and approximately $99\%$ validity for fewer than $5\cdot 10^{4}$ simulations, TSNPE-PF outperforms all comparison methods, which either required substantially larger simulation budgets to reach comparable performance, or did not reach it at all.
Our results demonstrate that TSNPE-PF can infer posteriors with substantially improved simulation efficiency, making it a powerful approach for complex, high-dimensional models with a limited simulation budget.
@@ 4 Discussion
We have shown here that pre-trained foundation models—specifically, TabPFN—can be effectively repurposed to perform accurate training-free simulation-based inference. Together with autoregressive sampling for high-dimensional posterior distributions, filtering to increase context size, and fast density evaluation using the density ratio, NPE-PF provides a flexible inference method that can be effortlessly applied to many different inference tasks. Across several benchmarks and two real-world tasks, NPE-PF consistently estimates accurate posteriors with particularly strong results for small simulation budgets, often requiring orders of magnitude fewer simulations than existing SBI methods. These results establish foundation models as a promising direction for training-free inference in scientific applications.
Related work.
@@
Our work relates to meta-learning [64], where models generalize across tasks by acquiring shared inductive biases. Meta learning has been applied to amortized (Bayesian) inference [65, 66]. More specifically, however, NPE-PF is an instance of in-context learning (ICL), where task adaptation is achieved solely by conditioning pre-trained models on different inputs [67]. ICL was popularized by LLMs [68], but its scope has expanded beyond language, with work showing that transformers can learn various function classes in context [69, 70] and work introducing prior fitted networks [71, 40, 35]. NPE-PF extends this direction by enabling ICL for simulation-based Bayesian inference. We note that this does not require training the inference networks on any SBI tasks, but rather works well using an “off-the-shelf” model trained for classification and regression.
@@
In addition to NPE-PF, multiple methods have been proposed to perform ICL for Bayesian inference [72, 73, 74]. These approaches parameterize posteriors using, for example, Gaussians, flows, or mixtures and require fixed input and parameter dimensionality and, to date, are only trained on a specific, limited class of probabilistic models. In contrast, NPE-PF leverages autoregressive modeling over parameter dimensions, enabling zero-shot generalization without retraining across inference problems with varying shapes.
@@
Finally, several works have proposed extensions for TabPFN beyond its original setting, targeting new data modalities, larger context sizes, or higher-dimensional features. To address TabPFN’s limited context size, retrieval-based approaches select informative subsets of the training data [42, 75, 44, 43], including $k$-NN-based filtering [42, 44, 43], which we adapt in NPE-PF for Bayesian inference. Crucially, we demonstrate that this filtering does not bias the posterior, as the filter step can be interpreted as a calibration kernel [5]. Other strategies include data distillation [76, 77], boosting weak learners [78], feature-wise ensembling [79], encoders and bagging [46], as well as non-autoregressive extensions to time series forecasting [80] and tabular data generation [81]. TabICL [82] proposes architectural changes to scale to datasets with over $10^{5}$ points. NPE-PF departs from these approaches by repurposing TabPFN for posterior inference through filtering and fast evaluation of densities.
@@
Limitations. While NPE-PF eliminates the need for simulation-specific training, it inherits properties of TabPFN. TabPFN has a soft limit in maximum feature size ($\leq 500$) and context length ($\leq 10^{4}$). We present methods to overcome these limitations, i.e, filtering and embedding nets. However, as the number of simulations and the need for density evaluations increase, the performance advantage of NPE-PF over standard methods such as NPE diminishes. Furthermore, while embedding networks can increase the available feature size, they require separate pre-training. In addition, modeling complex parameter spaces—such as spatially varying or continuous parameters—remains challenging. Finally, like other in-context methods, NPE-PF is more expensive than flow-based NPE, a challenge exacerbated by the autoregressive estimation of posteriors. However, unlike other in-context methods for Bayesian inference, NPE-PF does not require training and hyperparameter tuning, making it suitable for exploratory workflows or non-expert users. In addition, by requiring substantially fewer simulations, NPE-PF accelerates the overall SBI workflow. Thus, while we expect NPE-PF to be particularly useful for exploratory workflows or expensive simulators, inference methods based on trained density estimators might still be advantageous in settings where large numbers of simulation are available, or where bespoke embedding networks and fast sampling are required, such as in some applications in astrophysics [83].
Conclusion.
@@
We here presented NPE-PF, a method that repurposes the tabular foundation model TabPFN for training-free simulation-based Bayesian inference. NPE-PF is competitive with—and often outperforms—existing methods, especially for small simulation budgets. NPE-PF is a simple-to-use and simulation-efficient tool for simulation-based inference, with the potential to become the go-to method for novice users, exploratory workflows, and applications with constrained simulation budgets.

### v2

@@ 3 Experiments
To assess NPE-PFN, we conduct experiments on synthetic SBI benchmark tasks and real data, covering scenarios from low to high-dimensional data and including cases with model misspecification.
Figure 2: SBI benchmark results for amortized and sequential NPE-PFN. (a) C2ST for NPE, NLE, and NPE-PFN across ten reference posteriors (lower is better); dots indicate averages and bars show 95% confidence intervals over five independent runs. (b) Average time to generate $10^{4}$ posterior samples and (if applicable) training time. (c) C2ST for sequential methods.
@@ 3.1 Amortized and sequential benchmark performance
We evaluate NPE-PFN on various tasks from the SBI benchmark [27], which provides ground truth posterior samples for 10 observations for each task. We measure posterior sample quality using the classifier two-sample test (C2ST, 53). A C2ST accuracy of 0.5 indicates that the approximate posterior exactly matches the ground truth posterior, as the classifier fails to distinguish between the two sample sets. In contrast, an accuracy of 1.0 reflects a strong mismatch between the estimated and true posterior. As baselines, we compare against flow-based NPE and NLE provided by the SBI library [54, 55]. For sequential baselines, we use truncated sequential NPE (TSNPE [37]) and sequential NLE (SNLE [11]). Posterior estimates are computed across four simulation budgets, ranging from $10^{2}$ to $10^{5}$ simulations. Beyond $10^{4}$ simulations, NPE-PFN incorporates filtering, whereas an alternative unfiltered variant, denoted NPE-PFN (unfiltered), extends the context size beyond the TabPFN-recommended limit. For budgets $\leq 10^{4}$, both variants are identical.
@@
Amortized inference. For $10^{2}$ simulations, NPE-PFN substantially outperforms NLE and NPE on all but one task (Fig. 2a,b). For $10^{3}$ or $10^{4}$ simulations, NPE-PFN is generally competitive, with the exception of a few tasks where some baselines outperform it. For example, NLE outperforms NPE-PFN on the simple-likelihood-complex-posterior (SLCP) task, as NPE-PFN, like NPE, directly targets the posterior instead of the simpler likelihood. However, NPE-PFN performs substantially better than both baselines on tasks such as SIR or Lotka-Volterra for all simulation budgets. With $10^{5}$ simulations, the difference between NPE-PFN and its unfiltered variant becomes evident. The extended context does not improve performance for larger simulation budgets, while the filtered NPE-PFN variant achieves significantly better C2ST accuracies, matching or surpassing the baselines. While NPE-PFN is training-free, inference speed depends on the number of simulations, and the dimensionality of the parameter and observation spaces. For the benchmark tasks, we observe an inference speed that is comparable to NLE (Fig. 2b), but orders of magnitude slower than NPE, which is near-instantaneous. We perform a careful analysis of the inference speed of NPE-PFN across all relevant variables for users to make an informed decision in their respective application (Appendix Sec. D.3, Tab. D-2).
The advantage of NPE-PFN over baseline methods remains for other tasks outside the SBI benchmark suite (Appendix Sec. D.1, Fig. D-1), as well as compared to other baseline methods such as NRE, NPE ensembles (Appendix Sec. D.2, Fig. D-2), and even when performing extensive hyperparameter optimization for NPE (Appendix Fig. D-4). Furthermore, NPE-PFN posterior estimates are well-calibrated (Fig. D-3).
In cases where data dimension $d_{x}$ outgrows the supported feature dimension and learned summary statistics are required, the end-to-end trained NPE, however, outperforms NPE-PFN with independently pre-trained summary statistics (Appendix Sec. D.4, Fig. D-5). In addition, we ablate the effect of the filter size indicating that the default size is (close to) optimal for all considered tasks (Appendix Sec. D.6, Fig. D-7). Finally, we vary the autoregressive order, in which the parameter dimensions are sampled, and find no noticeable effect compared to the default order (Appendix Sec. D.8).
@@
Sequential inference. We adapt the setup in Deistler et al. [37] to evaluate TSNPE-PFN. Rejection sampling is performed using the fast, approximate ratio-based densities (Sec. 2.4, Appendix Sec. D.5, Fig. D-6 for a comparison between autoregressive and ratio-based density evaluation). We equally divide the simulation budget into 10 rounds, with proposals truncated to the $1-\varepsilon$ highest density region, where $\varepsilon=10^{-3}$. TSNPE-PFN excels with small simulation budgets (Fig. 2c). On the SIR task, it reaches close to optimal C2ST accuracy with only $10^{2}$ simulations. At larger budgets, TSNPE-PFN maintains a strong performance, particularly on the Lokta-Volterra task, which is highly challenging for other SBI methods. TSNPE-PFN benefits from NPE-PFN’s ability to generate high-quality posterior estimates with few simulations, allowing it to acquire more relevant simulations in early rounds.
@@
Figure 3: Robustness under misspecification. C2ST against a well-specified ground truth posterior. The red star marks the well-specified model. Rows: Prior and likelihood misspecification. Columns: NPE, NPE-PFN, and their difference in terms of C2ST accuracy (darker blue indicates that NPE-PFN is better).
Summary. NPE-PFN provides accurate posterior estimates, especially for small simulation budgets. With filtering, NPE-PFN is also competitive for larger simulation budgets. As an in-context learning method, it requires no training on specific simulations and offers inference speeds comparable to other baselines, though runtime scales with the number of simulations and the dimensionality of the parameters and observations. This flexibility makes NPE-PFN applicable to a wide range of setups and simulators. We observe NPE-PFN performs particularly well in tasks with underlying graphical structure and conditional (in-)dependencies, which are similar to the structural causal models used in TabPFN’s pre-training. In addition to posterior estimation with NPE-PFN, we evaluate the unconditional high-dimensional density estimation capabilities of TabPFN on UCI datasets [56], showing that it is more data-efficient than a neural spline flow (Appendix Sec. D.9, Fig. D-10).
@@ 3.2 Impact of misspecification
To investigate how NPE-PFN handles misspecification, we use the 2D Gaussian means misspecification benchmark from Schmitt et al. [57], considering both prior and likelihood misspecification. The ground truth is given by the prior ${\bm{\mu}}\sim\mathcal{N}({\bm{0}},{\bm{I}})$ and likelihood ${\bm{x}}_{i}\sim\mathcal{N}({\bm{\mu}},{\bm{I}})$. Prior misspecification uses the prior ${\bm{\mu}}\sim\mathcal{N}({\bm{\mu}}_{m},\tau_{m}{\bm{I}})$ while keeping the true likelihood; Likelihood misspecification uses the true prior with the likelihood ${\bm{x}}_{i}\sim\lambda_{m}\text{Beta}(2,5)+(1-\lambda_{m})\mathcal{N}({\bm{\mu}},\tau_{m}{\bm{I}})$, where $\tau_{m}\in\mathbb{R}^{+}$, $\lambda_{m}\in[0,1]$. We evaluate the C2ST of predicted posteriors against the reference posterior distribution of the well-specified data using $10^{2}$ simulations (details in Appendix Sec. 3.2).
Both NPE-PFN and NPE degrade as the level of misspecification increases, with prior misspecification leading to more severe performance degradation than likelihood misspecification (Fig. 3). However, NPE-PFN achieves better C2ST values than NPE across a broader range of misspecification levels. In addition, NPE-PFN remains insensitive to various feature and noise distributions (Appendix Sec. D.7, Fig. D-8). Finally, we find that the performance of NPE-PFN on the SBI benchmark tasks is invariant to the choice of the autoregressive order (Appendix Sec. D.8, Fig. D-9).
Thus, the trend observed in the SBI benchmark extends to these simple but misspecified tasks, on which NPE-PFN also consistently outperforms NPE. Next, we evaluated whether this performance gap also holds for scientific tasks with real data, on which the simulator is potentially misspecified.
Figure 4: Posterior inference for observations from the Allen cell type database. (a) Average distance to observation in standardized space of summary statistics for both the real and synthetic observations. (b) Posterior predictive simulations for real observations using TSNPE-PFN and TSNPE. (c) Posterior marginals for one synthetic observation; TSNPE-PFN marginals are substantially more constrained for several parameters. (d) Simulation-based calibration for NPE-PFN and NPE on the $H$ simulator.
@@ 3.3 Single-Compartment Hodgkin-Huxley Model
We evaluate the performance of (TS)NPE-PFN on a challenging inference task, considering the Hodgkin-Huxley (HH) simulator of single-neural voltage dynamics [58, 59]. Following Gonçalves et al. [60], the simulator has eight parameters and seven summary statistics based on the voltage trace. We infer posteriors for 10 real observations from the Allen cell type database [61], for which inference is known to be challenging [62, 63, 64]. We evaluate inference quality via posterior predictives and the average Euclidean distance between predictives and real observations in a standardized summary statistics space. In addition, for each real observation, we create an associated synthetic, and therefore well-specified observation, where the ground truth parameter is known, allowing for direct comparison between inferred posteriors and true values.
TSNPE-PFN yields posterior predictives that closely match real observations with a simulation budget of $10^{4}$ (Fig. 4a,b). It outperforms TSNPE, which struggles to produce realistic predictives, especially at smaller budgets. For synthetic observations, TSNPE-PFN again achieves better predictive accuracy with fewer simulations, with the average distance approaching zero for a simulation budget of $10^{4}$. In parameter space, TSNPE-PFN posteriors are tightly concentrated around the true parameters, with significant improvement over TSNPE (Fig. 4c). In the amortized setting, NPE-PFN shows good calibration for $10^{3}$ and $10^{5}$ simulations but is slightly overconfident for $10^{4}$ (Fig. 4d). However, (TS)NPE-PFN achieves significantly better prediction quality and yields tight posteriors around the true parameters, making it more sensitive to miscalibration.
These results demonstrate TSNPE-PFN’s superior simulation efficiency, achieving the same predictive quality as TSNPE with an order of magnitude fewer simulations. This efficiency makes TSNPE-PFN particularly well-suited for inference with real, potentially misspecified simulators under limited simulation budgets.
@@ 3.4 Pyloric Network Model
Finally, we evaluate TSNPE-PFN on a high-dimensional simulator of the pyloric network in the stomatogastric ganglion (STG) of the crab Cancer Borealis, a well-studied circuit generating rhythmic activity [65, 66, 60]. This model consists of three neurons, each with eight membrane conductances, interconnected by seven synapses, resulting in a 31-dimensional parameter space. The simulated voltage traces are summarized using 15 established statistics. A key challenge in this setting is the extreme sparsity of valid simulations. Specifically, 99% of parameter samples from the prior yield implausible voltage traces, preventing the computation of summary statistics [67].
For this reason, earlier work relied on millions of simulations for an amortized NPE-based posterior approximation (18 million in Gonçalves et al. [60]). The simulation count could later be reduced by restricting the prior to valid simulation using an additional classifier [67]. In addition, several sequential algorithms have been developed to address this problem [18, 37, 68], the latest of which achieves good posterior predictives using $1.5\cdot 10^{5}$ simulations.
Figure 5: Results on the pyloric simulator. (a) Voltage traces from the experimental measurement (top) and a posterior predictive simulated using the posterior mean from TSNPE-PFN as the parameter (bottom). (b) Average distance (energy scoring rule) to observation and percentage of valid simulation from posterior samples; compared to experimental results obtained in Glaser et al. [68].
To perform sequential inference on the pyloric network with TSNPE-PFN, we use a restricted prior based on the TabPFN classifier. Due to the high dimensionality of the parameter space, rejection sampling is prohibitive. Therefore, we use sampling importance resampling with an oversampling factor of 10. In the first round, we use $5\cdot 10^{3}$ simulations and run TSNPE-PFN for 45 rounds, adding $10^{3}$ simulations per round for a total of $5\cdot 10^{4}$ simulations (details in Appendix Sec. C.4).
The generated voltage traces from posterior mean estimates of TSNPE-PFN closely resemble true measurements (Fig. 5). With more than $50\%$ valid simulations at just $5\cdot 10^{3}$ simulations and approximately $99\%$ validity for fewer than $5\cdot 10^{4}$ simulations, TSNPE-PFN outperforms all comparison methods, which either required substantially larger simulation budgets to reach comparable performance or did not reach it at all. Furthermore, the autoregressive evaluation order again has no noticeable effect on the resulting performance (Appendix Sec. C.4).
Our results demonstrate that TSNPE-PFN can infer posteriors with substantially improved simulation efficiency, making it a powerful approach for complex high-dimensional models with a limited simulation budget.
@@ 4 Discussion
We have shown here that pre-trained foundation models—specifically, TabPFN—can be effectively repurposed to perform accurate training-free simulation-based inference. Together with autoregressive sampling for high-dimensional posterior distributions, filtering to increase context size, and fast density evaluation using the density ratio, NPE-PFN provides a flexible inference method that can be effortlessly applied to many different inference tasks. Across several benchmarks and two real-world tasks, NPE-PFN consistently estimates accurate posteriors with particularly strong results for small simulation budgets, often requiring orders of magnitude fewer simulations than existing SBI methods. These results establish foundation models and in-context learning as a promising direction for training-free inference in scientific applications.
@@
Related work. Our work relates to meta-learning [69], where models generalize across tasks by acquiring shared inductive biases. Meta learning has been applied to amortized (Bayesian) inference [70, 71]. More specifically, however, NPE-PFN is an instance of in-context learning (ICL), where task adaptation is achieved solely by conditioning pre-trained models on different inputs [43]. ICL was popularized by LLMs [72], but its scope has expanded beyond language, with work showing that transformers can learn various function classes in context [73, 74] and work introducing prior fitted networks [44, 45, 38]. NPE-PFN extends this direction by enabling ICL for simulation-based Bayesian inference. We note that this does not require training the inference networks on any SBI tasks, but rather works well using an “off-the-shelf” model trained for classification and regression.
@@
In addition to NPE-PFN, multiple methods have been proposed to perform ICL for Bayesian inference [75, 76, 77]. These approaches parameterize posteriors using, for example, Gaussians, flows, or mixtures and require fixed input and parameter dimensionality and, to date, are only trained on a specific, limited class of probabilistic models. In contrast, NPE-PFN leverages autoregressive modeling over parameter dimensions, enabling zero-shot generalization without retraining across inference problems with varying shapes.
@@
Finally, several works have proposed extensions for TabPFN beyond its original setting, targeting new data modalities, larger context sizes, or higher-dimensional features. To address TabPFN’s limited context size, retrieval-based approaches select informative subsets of the training data [47, 78, 49, 48], including nearest-neighbor-based filtering [47, 49, 48], which we adapt in NPE-PFN for Bayesian inference. Crucially, we demonstrate that this filtering does not bias the posterior, as the filter step can be interpreted as a calibration kernel [5]. Other strategies include data distillation [79, 80], boosting weak learners [81], feature-wise ensembling [82], encoders and bagging [51], as well as non-autoregressive extensions to time series forecasting [83] and tabular data generation [84]. TabICL [85] proposes architectural changes to scale to datasets with over $10^{5}$ points. NPE-PFN departs from these approaches by repurposing TabPFN for posterior inference through filtering and fast evaluation of densities.
@@
Limitations. While NPE-PFN eliminates the need for simulation-specific training, it inherits properties of TabPFN. TabPFN has a soft limit in maximum feature size ($\leq 500$) and context length ($\leq 10^{4}$). We present methods to overcome these limitations, i.e, filtering and embedding nets. However, as the number of simulations and the need for density evaluations increase, the performance advantage of NPE-PFN over standard methods such as NPE diminishes. Furthermore, while embedding networks allow NPE-PFN to be used for high-dimensional observations, they require separate pre-training. In addition, modeling high-dimensional parameter spaces beyond TabPFN’s maximum feature size of 500 dimensions remains challenging. Finally, like other in-context methods, NPE-PFN is more expensive than flow-based NPE, a challenge exacerbated by the autoregressive estimation of posteriors. However, unlike other in-context methods for Bayesian inference, NPE-PFN does not require training and hyperparameter tuning, making it suitable for exploratory workflows or non-expert users. In addition, by requiring substantially fewer simulations, NPE-PFN accelerates the overall SBI workflow. Thus, while we expect NPE-PFN to be particularly useful for exploratory workflows or expensive simulators, inference methods based on trained density estimators might still be advantageous in settings where large numbers of simulations are available, or where bespoke embedding networks and fast sampling are required, such as in some applications in astrophysics [22]. An interesting direction for future work is to transfer NPE-PFN’s advantages to existing, faster inference techniques (e.g., via student–teacher distillation).
@@
Conclusion. We presented NPE-PFN, a method that repurposes the tabular foundation model TabPFN for training-free simulation-based Bayesian inference. NPE-PFN is competitive with—and often outperforms—existing methods, especially for small simulation budgets. NPE-PFN is a simple-to-use and simulation-efficient tool for SBI, with the potential to become the go-to method for novice users, exploratory workflows, and applications with constrained simulation budgets.